An older person is 7 years older than four times the age of a younger person. The sum of their ages is 17. Find their ages.
Answer by addingup(3677)

Older: o
Younger: y
----------------
o + y = 17 subtract y from both sides:
o = 17 - y
--------------------
o = 4y + 7 substitute for o in the original equation above:
4y + 7 + y = 17
5y = 10
y = 2
o = 17 - y = 17 - 2 = 15
-------------------------------
Check:
O = 4y + 7
o = 4(2) + 7
o = 8 + 7 = 15 correct.